A gas filled a 5 dm3 container at a pressure of 3.0 atm, at 301 K. If the pressure of the system is doubled, what is the new volume that the gas will occupy at 301 K?


1
Expert's answer
2021-06-14T13:35:09-0400

P1 = 3.0 atm

V1 = 5 dm3

P2 = (3.0 * 2) = 6.0 atm

V2 = unknown

T1 = T2 = 301 K = const

 

Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used. 

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the new volume, solve the equation for V2:

V2 = P1V1 / P2

V2 = (3.0 atm × 5 dm3) / (6.0 atm) = 2.5 dm3

V2 = 2.5 dm3


Answer: 2.5 dm3 is the new volume if gas.